As part of an interview for a summer job with the Coast Guard, you are asked to help determine the search area for two sunken ships by calculating their velocity just after they collided. According to the last radio transmission from the 40,000-ton luxury liner, the Hedonist, it was going due west at a speed of 20 knots in calm seas through a rare fog just before it was struck broadside by the 60,000-ton freighter, the Ironhorse, which was traveling north at 10 knots. The transmission also noted that when the freighter's bow pierced the hull of the liner, the two ships stuck together and sank together.
Answer #1

I have used the principle of Conservation of momentum to solve this question. The collision in this case is purely inelastic, as both the ships are stuck together after colliding with each other, hence their final momentum is calculated as a single object.

solution - We can solue this by applying the principle of consumation of momentum. In a collision, the total momentum is Сици

In case of the Freighter le = mu m = 60,000 ton W = 10 (j) knoti = my = 600,000 (ſ) ton thats Now, Total momentum before thespeed after collision = (40)2 + (30 knots speed = 7.143 knate tano = y = 30/7 =-3 - 40/7 o= tan- (-34) 0 = -370 Their direct